At a constant amount and volume an unknown gas has an initial pressure of 0.70 ATM at a 10.5°C. Calculate the pressure of the gas if the temperature increases to 37.0°C
Answer #1

According to gas laws pressure and temperature are related as,

P1/T1 = P2/T2

P1 & P2 = pressure. T1 & T2 = temperature

P1 = 0.70 atm. T1 = 10.5° C

P2 = ? T2 = 37° C

0.70/10.5 = P2/37

0.067 = P2/37

P2 = 0.067 × 37

P2 = 2.48 atm

Answer = 2.48 atm